OpenStudy (anonymous):

Solve for x: −5|x + 1| = 10 x = 0 x = −3 and x = 1 x = −1 and x = 3 No solution

OpenStudy (welshfella):

Hint by definition |x + 1| is positive

OpenStudy (anonymous):

these confuse me

OpenStudy (welshfella):

−5|x + 1| = 10 |x + 1| = 10/-5 = -2 so?

OpenStudy (welshfella):

the vertical lines mean the absolute value of x+1 which must be positive

OpenStudy (welshfella):

absolute value is the positive value

OpenStudy (anonymous):

so would it be x= -3 x= 1 ?

OpenStudy (welshfella):

No as I said |x + 1| is positive

OpenStudy (welshfella):

check my second post

OpenStudy (anonymous):

i dont see how to solve that

OpenStudy (anonymous):

10/-5 = -2

OpenStudy (welshfella):

what value did i get for |x + 1| in my second post?

OpenStudy (welshfella):

|x + 1| = 10/-5 = -2

OpenStudy (anonymous):

-2 ?

OpenStudy (welshfella):

yes - now does that make sense given what i said about the sign of |x+1|?

OpenStudy (anonymous):

kind of yeah.

OpenStudy (welshfella):

|x + 1| must be positive but we got it equal to -2 so? Look at the choices.

OpenStudy (anonymous):

C ? im really bad with this sorry

OpenStudy (anonymous):

no solution ? since you we got it to equal -2

OpenStudy (welshfella):

No its d , because -2 is negative and |x + 1| must be positive there is no solution.

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

the vertical lines make whats inside them positive for example |-5| = 5

OpenStudy (anonymous):

you help a lot. i understand it a lot better

OpenStudy (welshfella):

yw